OpenStudy (anonymous):

I need help with the follow equation: (Please explain how you got to the answer.)

OpenStudy (anonymous):

\[\frac{ 160 x^{2} }{ 20 }\div \frac{ 4 }{ 5y ^{2} }\]

OpenStudy (anonymous):

Are you just looking to simplify this? It's not really an equation(no equal sign), so there's nothing to solve for. To simplify here i would simply combine the two fractions. Dividing a fraction is the same as flipping it over and multiplying so: \[160x^2/20 \div 4/5y^2 = 160x^2 * 5y^2 / 20 *4\] 160*5 on the top and 80 on the bottom works out to just 10 on the top: \[10x^2y^2\]

OpenStudy (anonymous):

Okay so I just need to flip any fraction like that in which I am dividing to get my answer and then cross multiply?

OpenStudy (anonymous):

Yeah, if you're dividing one fraction by another just flip the second one and multiply. Cross multiplying usually refers when you've got one fraction equal to another one so if you have \[x/350 = 90/100\]Cross multiplying turns it into: \[90*350 = 100x\]\[315 = x\]Like if you wanted to find out how many points out of 350 would give you 90%(90/100)

OpenStudy (anonymous):

So does cross multiply work for this sort of problem?

OpenStudy (anonymous):

Oh and this is a "equals" problem. Forgot to add the sign.

OpenStudy (anonymous):

Assuming that's supposed to be an = in the center instead of the division sign, then yeah you would cross multiply here. It'ld\[ y = \pm \sqrt{1/10x^2}, x = \pm \sqrt{1/10y^2}\]be \[80 = 160*5x^2y^2\]\[ 1 = 10x^2y^2\]Solve for whatever: \[y = \pm \sqrt{1/10x^2}\] The graph ends up looking like a big diamond.

OpenStudy (anonymous):

This is what the problem looks like on MyMathLab. Whatever you did up there seem to be way over my head. I am in College Intermediate Algebra...

OpenStudy (anonymous):

Ah, kk, they only wanted you to simplify. The first answer is pretty much what they were looking for then.

OpenStudy (anonymous):

I tried it and that did not work. It is supposed to still be in a fraction form?

OpenStudy (anonymous):

The denominator turns into 1 when you simplify so there's no fraction to write. You do have x^2 in your OP and x^3 on the question you just pasted over, but outside of that there's really not much else going on.

OpenStudy (anonymous):

Okay so is the x only x and not x^2?

OpenStudy (anonymous):

Never mind I got it... I don't know what I was thinking. Thank you for your patients with me and your help!